OpenStudy (anonymous):

At 7:00 A.M. two planes leave an airport, one flying south at 475km/h, the other flying north at 650km/h. At what time will they be 4500 km apart?

OpenStudy (anonymous):

What is their relative velocity? duration - (distance)/(relative velocity) Time = 7 a.m. + duration.

OpenStudy (anonymous):

plzz help.

OpenStudy (anonymous):

If two objects are going in opposite directions, their relative velocity is the sum of their speeds, here (650 + 475) = 1125 km/h

OpenStudy (anonymous):

Since they're flying away from each other, the distance between them is increasing by the sum of their speeds, which is 475 + 650 =1025 km/h. To find the time it takes for the distance between them to be 4500km, divide this distance by the speed.\[\Delta t=\frac{\Delta x}{v}=\frac{4500km}{1025km/hr}\approx4.39hr\]Now that we know how much time has passed, just add it to the starting time to get the final time.\[t_f=t_i+\Delta t=7hr+4.39hr=11.39hr\]11.39 doesn't make sense as a time so convert .39hr to minutes to get the final answer: 11:23AM

OpenStudy (anonymous):

Thanks a lot, that made much more sense:)

OpenStudy (anonymous):

Good explanation, although 650+475 = 1125 on my calculator.

OpenStudy (anonymous):

Oh thanks, so change in time would be exactly 4 hours so the end time would be 11:00AM, that's much cleaner.

OpenStudy (anonymous):

You are very welcome. Cleaner is nicer. ["beauty is truth..."]